OpenStudy (anonymous):

Find all points (x,y) on the graph of f(x)=2x^2-3x with tangent lines parallel to the line y=9x+4

OpenStudy (michele_laino):

we have to compute the first derivative of your function, first

OpenStudy (michele_laino):

do you know how to compute such first derivative?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

it would be 4x-3

OpenStudy (michele_laino):

ok! now using the geometrical meaning of the first derivative, we have to solve this equation: \(4x-3=9\) being \(9\) the slope of the line

OpenStudy (michele_laino):

please solve that equation with respect to \(x\) what do you get?

OpenStudy (anonymous):

x=3

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

we got only one x-coordinate, so the y-coordinate is: \(y=2\cdot 3^2-3\cdot 3=...\)

OpenStudy (anonymous):

y=9

OpenStudy (michele_laino):

correct! So the requested point is \((3,9)\)

OpenStudy (anonymous):

It says all points is (3,9) the only one?

OpenStudy (michele_laino):

yes! It is the only point, since the first derivative is a polynomial of degree 1

OpenStudy (anonymous):

Ok thank you very much

OpenStudy (michele_laino):

:)